The most recent financial statements for Bradley, Inc., are shown here (assuming no income taxes):

Income Statement
Sales $ 9,000
Costs (6,750 )
Net income $ 2,250
Balance Sheet
Assets $ 28,800 Debt $ 10,200
Equity 18,600
Total $ 28,800 Total $ 28,800

Assets and costs are proportional to sales. Debt and equity are not. No dividends are paid. Next year’s sales are projected to be $10,350. What is the external financing needed? (A negative value should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to the nearest whole number.)
Answer #1

Projected sale is $10,350 which is $1350 [$10,350-$9,000] more than recent year sales of $9,000

%change in sales = (change in sales / recent year sales)×100

%change in sales = ($1350/$9000)×100=15%

pro-forma for projected income statement

Particulars current year($) Projected($)
Sales 9,000 10,350
Costs (6750) (7,762.5)...[working note 1]
Net Income 2250 2,587.5

Working note 1- costs change according to sales

Here in this case sales changes with 15%,so cost would also increase by 15%

$6,750×15% = $6,750×15/100= $1,012.5

Estimated Cost would be $6,750+$1,012.5 =$7,762.5

pro-forma of balance sheet

Assets Current Year($) Projected($) Labilities Current Year($) Projected($)
Assets 28,800 33,120...(working note 2) Debt 10,200 10,200
Equity 18,600 21,187.5....(working note 3)
External financing 1,732.5(working note 4)
28,800 33,120 28,800 33,120

Working Notes:

2) Projected assets is to be changed according to sales proportion(15%) calculated in working note 1

So 15%change in assets = $28,800×15% =$28,800×15/100=$4,320

Projected assets = $28,800+$4,320=$33,120

3) New Equity = beginning equity + projected net income + Dividend

New Equity =$18,600+$2,587.5+0

New Equity =$21,187.5

4) External financing needed

Projected Assets = projected Liabilities

$33,120= $10,200+$21,187.5+ external finance

External finance = $33,120-($10,200+$21,187.5)

External financing = $33,120- $31,387.5= $1732.5

hence External financing needed is $1,732.5